# Detecting Out-of-distribution through the Lens of Neural Collapse

### v5

@@ 3.3 Out-of-Distribution Detection
Based on our understanding, we design an efficient and versatile OOD detector. Specifically, we propose to detect OOD based on feature proximity to the weight vectors of the predicted class. For the proximity metric, we avoid Euclidean-based metrics as they require estimating the scaling factor $\lambda$ in Equation 2. This estimation tends to be imprecise for general classifiers which may cease training prior to convergence, resulting in suboptimal performance of Euclidean-based metrics shown in Appendix B. Instead, we design an angle-based metric, adjusted for class-wise difference. Specifically, we propose to quantify the proximity as the norm of projection of the weight vector $\bm{w}_{c}$ onto the centered feature $\bm{h}-\bm{\mu}_{G}$, where $c$ corresponds to the predicted class, i.e.,
@@
where $cos(\bm{w}_{c},\bm{h}-\bm{\mu}_{G})=\frac{(\bm{h}-\bm{\mu}_{G})\cdot\bm{w}_{c}}{\|\bm{h}-\bm{\mu}_{G}\|_{2}\|\bm{w}_{c}\|_{2}}$. A higher $\mathtt{pScore}$ indicates closer proximity to the weight vector and thus a lower chance of OOD-ness. Geometrically, thresholding on $\mathtt{pScore}$ selects infinite hyper-cones centered at the weight vectors, as illustrated in Figure 1 Right. Within the same predicted class, $\mathtt{pScore}$ is proportional to the cosine similarity. Across different classes, $\mathtt{pScore}$ adapts to class-wise difference by selecting wider hyper-cones for classes with larger weight vectors, which tend to have larger decision regions. As shown in Appendix B, our $\mathtt{pScore}$ with class-wise adjustment outperforms vanilla cosine similarity. Notably, our $\mathtt{pScore}$ incorporates class-specific information into characterizing ID clustering by using the weight vectors of the predicted class. This brings in additional gain in detection effectiveness, as we shall see in Section 4.
While $\mathtt{pScore}$ enhances efficiency and effectiveness, its performance is intrinsically contingent on the strength of ID clustering. Such contingency, widely exhibited by clustering-based methods (Lee et al., 2018; Sun et al., 2022; Tack et al., 2020), poses challenges on classifiers with less pronounced ID clustering, such as ImageNet ResNet-50 in Section 4.1. To mitigate such discrepancy, we complement $\mathtt{pScore}$ by considering the distance of ID clusters to the origin. Specifically, we enhance our proximity score by incorporating feature norms to filter out OOD near the origin, as illustrated in Figure 1 Right. Taking $\mathtt{L1}$ norm as an example, we define our detection score as $\mathtt{pScore}+\alpha\|\bm{h}\|_{1}$, where $\alpha$ controls the filtering strength and can be selected from a validation set as detailed in Section 4. We refer readers to Section 4.3 for the effect of different orders of $p$-norm. Thresholding on the detection score, we have Neural Collapse Inspired OOD Detector ($\mathtt{NCI}$): A lower score indicates a higher chance of OOD-ness.
$\mathtt{NCI}$ has $O(P)$ complexity, where $P$ represents the dimension of the penultimate layer. The complexity theoretically ensures computational scalability of $\mathtt{NCI}$ on large models. Empirically, $\mathtt{NCI}$ maintains inference latency comparable to the vanilla softmax-confidence detector, as we shall see in Section 4.
@@ 4 Experiments
Table 1: $\mathtt{NCI}$ achieves high AUROC, low FPR95, and low latency on CIFAR-10 and ImageNet OpenOOD benchmarks. The CIFAR-10 classifier is a pre-trained ResNet-18, and the ImageNet classifier is a pre-trained ResNet-50.
@@
(a) $\mathtt{NCI}$ remains effective while baselines exhibit discrepancy. $\uparrow$ and $\downarrow$ denotes better performance. Bold indicates the best results while underline denotes the 2nd and 3rd best results. Methods with * are hyperparameter-free. Scores, except for the most recent baselines – $\mathtt{fDBD}$, $\mathtt{NECO}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – are copied from OpenOOD report (Zhang et al., 2023).
(b) $\mathtt{NCI}$ achieves simialr latency as $\mathtt{MSP}$, outperforming representative baselines. Per image latency reported.
Table 2: $\mathtt{NCI}$ achieves high AUROC, low FPR95 on ImageNet OpenOOD benchmarks across ViT B/16 and Swin v2 classifiers. Both transfomers are finetuned on ImageNet. While baselines exhibits discrepency, our $\mathtt{NCI}$ boosts Swin v2 performance while maintaining ViT effectiveness, even without filtering. Bold indicates the best results while underline denotes the 2nd best results.
@@
In this section, we extensively demonstrate the versatility and efficiency of $\mathtt{NCI}$ across classification tasks: CIFAR-10, CIFAR-100 (see App. D), ImageNet, as well as model architectures: ResNet, DenseNet (see App. D), ViT, Swin. We compare $\mathtt{NCI}$ against thirteen baseline methods, including the most recent ones, and demonstrate that $\mathtt{NCI}$ effectively mitigates the generalization discrepancies of existing methods. Following the OpenOOD benchmark (Zhang et al., 2023), we evaluate on six OOD sets for CIFAR-10 and CIFAR-100 classifiers and five for ImageNet classifiers. Performance is evaluated using two widely recognized metrics: the False Positive Rate at 95% True Positive Rate (FPR95) and the Area Under the Receiver Operating Characteristic Curve (AUROC). Lower FPR95 and higher AUROC values indicate better performance. We also report the per-image inference latency (in milliseconds) evaluated on a Tesla T4 GPU. In our experiments, other than the ablation study in Section 4.3, we use the $L1$-norm as the filtering term and select the filtering strength $\alpha$ from $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$ based on a validation set generated per pixel from Gaussian $N(0,1)$, following Sun et al. (2021) and Sun and Li (2022). In Section 4.3, we explore different choices of norm order and examine the effect of filtering on alternative clustering-based method. For detailed setups, please see Appendix A. We emphasize that our method and all baselines are post-hoc methods, while all models used are off-the-shelf and do not undergo prolonged training.
4.1 Versatility across Classification Tasks
@@
We first assess the performance of $\mathtt{NCI}$ and baselines across CIFAR-10 and ImageNet classification tasks. The two tasks provide an ideal test bed for evaluating versatility, as they drastically differ in input resolution, number of classes, and classification accuracy. We use the ResNets from the OpenOOD benchmark (Zhang et al., 2023): the CIFAR-10 classifier is a ResNet-18 with an accuracy of 95.06% and the ImageNet classifier is a ResNet-50 with an accuracy of 76.65%. Based on validation results, we set the filter strength $\alpha$ of the $L1$-norm to $10^{-2}$ for CIFAR-10 experiments and $10^{-3}$ for ImageNet experiments.
@@
Datasets For CIFAR-10 experiments, We follow the OpenOOD split of CIFAR-10 test set and evaluate on the OpenOOD benchmarks, including CIFAR-100 (Krizhevsky et al., 2009), Tiny ImageNet (Le and Yang, 2015), MNIST (Deng, 2012), SVHN (Netzer et al., 2011), Texture (Cimpoi et al., 2014), and Places365 (Zhou et al., 2017). For ImageNet experiments, we follow the OpenOOD split of ImageNet test set and evaluate on the OpenOOD benchmarks, including SSB-hard (Vaze et al., 2021), NINCO (Bitterwolf et al., 2023), iNaturalist (Van Horn et al., 2018), Texture (Cimpoi et al., 2014), and OpenImage-O (Wang et al., 2022).
@@
Baselines In Table 1(a), we compare our method with thirteen baselines. Some baselines focus more on the CIFAR-10 Benchmark while others focus more focused on the Imagenet Benchmark. Based on performance, we categorize the baselines, besides the vanilla confidence-based $\mathtt{MSP}$ (Hendrycks and Gimpel, 2016), into two groups: the “CIFAR-10 Strong" baselines, including $\mathtt{ODIN}$ (Liang et al., 2018), $\mathtt{Energy}$ (Liu et al., 2020), $\mathtt{Mahalanobis}$ (Lee et al., 2018), $\mathtt{KNN}$ (Sun et al., 2022), $\mathtt{ViM}$ (Wang et al., 2022), and $\mathtt{fDBD}$ (Liu and Qin, 2023); the “ImageNet Strong" baselines, including $\mathtt{GradNorm}$ (Huang et al., 2021), $\mathtt{NECO}$ (Ammar et al., 2023), $\mathtt{React}$ (Sun et al., 2021), $\mathtt{Dice}$ (Sun and Li, 2022), $\mathtt{ASH}$ (Djurisic et al., 2022), $\mathtt{Scale}$ (Xu et al., 2023). For details of the baselines, please see Appendix C.
OOD detection performance Table 1(a) highlights the discrepancy among existing methods: CIFAR-10 Strong baselines perform sub-optimally on ImageNet benchmarks, whereas ImageNet Strong baselines cannot achieve state-of-the-art performance on CIFAR-10 benchmarks. Conversely, our $\mathtt{NCI}$ largely mitigates the generalization gap, achieving competitive performance across both tasks. Moreover, $\mathtt{NCI}$ is practically as efficient as $\mathtt{MSP}$, as shown in Table 1(b) [^1] Running time of $\mathtt{KNN}$ and $\mathtt{MDS}$ on ImageNet are copied from Table 4 in Sun et al. (2022).. This enhances the efficiency of feature-space methods, aligning with our complexity analysis in Section 3 and Appendix C.
We highlight the following pairs of comparison:
  • $\mathtt{NCI}$ v.s. $\mathtt{NCI}$ w/o filter: The comparison validates the effectiveness of our norm-based filtering and highlights its importance in enhancing generalizability. On the CIFAR-10 classifier, strong ID clustering allows our method to achieve state-of-art performance without additional filtering. Conversely, on the ImageNet ResNet-50, ID clustering is less prominent (see Appendix E). There, norm-based filtering effectively complements the performance.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{KNN}$: Compared to $\mathtt{KNN}$, $\mathtt{NCI}$ avoids auxiliary models and significantly enhances efficiency as shown in Table 1(b). Notably, without filtering, our hyperparameter-free score still outperforms KNN with selected parameters on ImageNet benchmarks. The performance gain validates the benefit of incorporating class-specific information in our design.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{ASH}$ / $\mathtt{Scale}$: Compared to both baselines, our $\mathtt{NCI}$ achieves competitive performance on the ImageNet benchmark while significantly enhancing the performance on the CIFAR-10 benchmark, thereby largely improving versatility across classification tasks. Additionally, $\mathtt{ASH}$ and $\mathtt{Scale}$ introduce in a small delay on the ImageNet benchmark due to their cost of activation sorting per image. We expect the delay to increase and the latency gap to widen on classifiers with a larger activation dimension.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{NECO}$: Like our $\mathtt{NCI}$, $\mathtt{NECO}$ (Ammar et al., 2023) is motivated by Neural Collapse. Similar to our $\mathtt{NCI}$ with filtering, $\mathtt{NECO}$ incorporates distance to the origin using max-logit. However, $\mathtt{NECO}$, exclusively focuses on the subspace analysis of features and does not utilize the classification head. Such subspace operations require expensive matrix multiplication, resulting in a noticeable inference latency shown in Table 1(b). Conversely, our $\mathtt{NCI}$ explores the interplay between features and the classification head, and thus integrates class-specific information revealed by Neural Collapse. Compared to $\mathtt{NECO}$, our $\mathtt{NCI}$ improves in both efficiency and effectiveness.
4.2 Versatility across Model Architectures
@@
Now we assess the performance of $\mathtt{NCI}$ and baselines across different architectures. Particularly, we study two transformer-based models: ViT B/16 (Dosovitskiy et al., 2020) and Swin-v2 (Liu et al., 2022). Both transformers are finetuned on ImageNet, achieving an accuracy of 81.14% and 82.94% respectively. We follow the setup of the OpenOOD ImageNet Benchmark, detailed in Section 4.1. And based on validation results, we set the filter strength $\alpha$ of the $L1$ norm to $10^{-3}$ for both classifiers. In Table 2, we observe strong baselines on ViT – $\mathtt{KNN}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – exhibit discrepancy when applied to Swin v2, echoing the observations in Ammar et al. (2023). Conversely, our $\mathtt{NCI}$, even without filtering, improves baseline performance on Swin v2 while maintaining superior effectiveness on ViT. The addition of filtering further enhances overall generalizability. Our experiments on ResNet (Section 4.1) and DenseNet (Appendix D) further validate the versatility of $\mathtt{NCI}$ across model architectures.
@@ 4.3 Ablation on the Filtering Effect
Table 3: Ablation on filtering norm on ImageNet OpenOOD Benchmark with ResNet-50 backbone. AUROC score is reported (higher is better). Bold denotes the best result. Filtering with $\mathtt{L1}$ norm outperforms alternative choice of norms across OOD datasets.
In Table 3, we assess different orders of $p$-norm as the filtering term, compared to the $L1$ norm used so far. To ensure a fair comparison, we report the best performance from the filter strengths $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. The rest of the setup follows the ImageNet benchmarks in Section 4.1. As shown in Table  3, filtering with $L1$ norm achieves the best performance across OOD datasets, aligning with prior observations (Huang et al., 2021; Park et al., 2023). Meanwhile, we observe that in rare scenarios, e.g., a ResNet-18 on CIFAR-10, the $\mathtt{L1}$ norm cannot effectively characterize OOD’s proximity to the origin, leading to no extra performance gain compared to simply thresholding on $\mathtt{pScore}$. In these cases, our algorithm benefits from its ability to automatically select a low filter strength based on validation results, effectively disregarding the filtering term.
We further apply $\mathtt{L1}$-norm based filtering to $\mathtt{KNN}$ to see if this perspective can mitigate the discrepancy of clustering-based methods in general. In Table 4 222 Note that we report our run of KNN here to ensure a fair evaluation of the filtering effect. Our results are very similar to the OpenOOD results reported in Table 1(a), with only marginal differences. , we report the the best performance of $\mathtt{KNN}$ from filter strengths $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. We observe a significant performance gain from adding the filter, which further validates our understanding of ID clustering landscape from Neural Collapse. Note that our method outperforms the standalone $L1$ norm as well as $\mathtt{KNN}$, before and after filtering.

### v7

@@ 3.3 Out-of-Distribution Detection
Based on our understanding, we design an efficient and versatile OOD detector. Specifically, we propose to detect OOD based on feature proximity to the weight vectors of the predicted class. For the proximity metric, we avoid Euclidean-based metrics as they require estimating the scaling factor $\lambda$ in Equation 2. This estimation tends to be imprecise for general classifiers which may cease training prior to convergence, resulting in suboptimal performance of Euclidean-based metrics shown in Appendix 8. Instead, we design an angle-based metric, adjusted for class-wise difference. Specifically, we propose to quantify the proximity as the norm of projection of the weight vector $\bm{w}_{c}$ onto the centered feature $\bm{h}-\bm{\mu}_{G}$, where $c$ corresponds to the predicted class, i.e.,
@@
where $cos(\bm{w}_{c},\bm{h}-\bm{\mu}_{G})=\frac{(\bm{h}-\bm{\mu}_{G})\cdot\bm{w}_{c}}{\|\bm{h}-\bm{\mu}_{G}\|_{2}\|\bm{w}_{c}\|_{2}}$. A higher $\mathtt{pScore}$ indicates closer proximity to the weight vector and thus a lower chance of OOD-ness. Geometrically, thresholding on $\mathtt{pScore}$ selects infinite hyper-cones centered at the weight vectors, as illustrated in Figure 2 Right. Within the same predicted class, $\mathtt{pScore}$ is proportional to the cosine similarity. Across different classes, $\mathtt{pScore}$ adapts to class-wise difference by selecting wider hyper-cones for classes with larger weight vectors, which tend to have larger decision regions. As shown in Appendix 8, our $\mathtt{pScore}$ with class-wise adjustment outperforms vanilla cosine similarity. Notably, our $\mathtt{pScore}$ incorporates class-specific information into characterizing ID clustering by using the weight vectors of the predicted class. This brings in additional gain in detection effectiveness, as we shall see in Section 4.
While $\mathtt{pScore}$ enhances efficiency and effectiveness, its performance is intrinsically contingent on the strength of ID clustering. Such contingency, widely exhibited by clustering-based methods [20, 39, 40], poses challenges on classifiers with less pronounced ID clustering, such as ImageNet ResNet-50 in Section 4.1. To mitigate such discrepancy, we complement $\mathtt{pScore}$ by using the distance of ID clusters to the origin. Specifically, we enhance our proximity score by incorporating feature norms to filter out OOD near the origin, as illustrated in Figure 2 Right. Taking $\mathtt{L1}$ norm as an example, we define our detection score as $\mathtt{pScore}+\alpha\|\bm{h}\|_{1}$, where $\alpha$ controls the filtering strength and can be selected from a validation set as detailed in Section 4. We discuss the effect of different orders of $p$-norm in Section 4.3. Thresholding on the detection score, we have Neural Collapse Inspired OOD Detector ($\mathtt{NCI}$): A lower score indicates a higher chance of OOD-ness.
$\mathtt{NCI}$ has $O(P)$ complexity, where $P$ is the penultimate layer dimension. The complexity theoretically ensures computational scalability of $\mathtt{NCI}$ on large models. Empirically, $\mathtt{NCI}$ maintains inference latency comparable to the vanilla softmax-confidence detector, as shown in Section 4.
@@ 4 Experiments
Table 1: $\mathtt{NCI}$ reduces discrepencies and improves overall performance on CIFAR-10 and ImageNet benchmarks with minimal latency. CIFAR-10 uses ResNet-18 and ImageNet uses ResNet-50.
@@
(a) $\mathtt{NCI}$ ranks top-three in both benchmarks, while baselines show greater variability. $\uparrow$ and $\downarrow$ denotes better performance. Bold marks best, underline 2nd / 3rd. Methods with * are hyperparameter-free. Scores, except for the most recent baselines – $\mathtt{fDBD}$, $\mathtt{NECO}$, $\mathtt{ASH}$, $\mathtt{Scale}$ – are from OpenOOD [49].
(b) $\mathtt{NCI}$ improves the overall performance while reducing latency compared to strong baselines. AUROC averaged across CIFAR-10 and ImageNet benchmarks in Table 1(a), with per image latency reported.
Table 2: $\mathtt{NCI}$ reduces discrepencies and improves overall performance on ImageNet benchmarks across ViT B/16 and Swin v2 classifiers. Bold marks best, underline 2nd
(a) $\mathtt{NCI}$ boosts Swin v2 while maintaining ViT effectiveness compared to baselines, even without filtering.
(b) $\mathtt{NCI}$ improves the overall performance. AUROC averaged across two architectures in Table 2(a).
@@
In this section, we extensively evaluate $\mathtt{NCI}$ across classification tasks: CIFAR-10, CIFAR-100 (see App. 12), ImageNet, as well as model architectures: ResNet, DenseNet (see App. 12), ViT, Swin. We compare $\mathtt{NCI}$ against thirteen baseline methods. While $\mathtt{NCI}$ may not achieve the best performance on individual benchmarks, it mitigates the exisitng generalization discrepancies and achieves the best overall performance with minimal latency. Following the OpenOOD benchmark [49], we evaluate on six OOD sets for CIFAR-10 and CIFAR-100 classifiers and five for ImageNet classifiers. Performance is evaluated using two widely recognized metrics: the False Positive Rate at 95% True Positive Rate (FPR95) and the Area Under the Receiver Operating Characteristic Curve (AUROC). Lower FPR95 and higher AUROC values indicate better performance. We also report the per-image inference latency (in milliseconds) evaluated on a Tesla T4 GPU. In our experiments, other than the ablation study in Section 4.3, we use the $L1$-norm as the filtering term and select the filtering strength $\alpha$ from $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$ based on a validation set generated per pixel from Gaussian $N(0,1)$, following [38, 37]. Our method and all baselines are post-hoc methods, while all models used are off-the-shelf and do not require complete Neural Collapse convergence. For detailed setups and additional baselines, please see Appendix 7 and Appendix 11, respectively.
4.1 Mitigating Discrepencies across ID Datasets
@@
We first assess $\mathtt{NCI}$ and baselines across CIFAR-10 and ImageNet classification tasks. The two tasks provide an ideal test bed for evaluating versatility, as they drastically differ in input resolution, number of classes, and classification accuracy. We use ResNets from OpenOOD [49]: ResNet-18 for CIFAR-10 (95.06% accuracy) and ResNet-50 for ImageNet (76.65% accuracy). Based on validation results, we set the filter strength $\alpha$ of the $L1$-norm to $10^{-2}$ for CIFAR-10 experiments and $10^{-3}$ for ImageNet experiments.
@@
Datasets For CIFAR-10 experiments, We follow the OpenOOD split of ID test set and evaluate on the OpenOOD benchmarks, including CIFAR-100 [18], Tiny ImageNet [19], MNIST [6], SVHN [31], Texture [4], and Places365 [50]. For ImageNet experiments, we follow the OpenOOD split of ID test set and evaluate on the OpenOOD benchmarks, including SSB-hard [42], NINCO [2], iNaturalist [41], Texture [4], and OpenImage-O [43].
@@
Baselines In Table 1(a), we compare our method with thirteen baselines. Some baselines focus more on the CIFAR-10 Benchmark while others focus more focused on the Imagenet Benchmark. Therefore, we categorize the baselines, besides the vanilla confidence-based $\mathtt{MSP}$ [12], into two groups: the “CIFAR-10 Strong” baselines, including $\mathtt{ODIN}$ [21], $\mathtt{Energy}$ [24], $\mathtt{Mahalanobis}$ [20], $\mathtt{KNN}$[39], $\mathtt{ViM}$ [43], and $\mathtt{fDBD}$ [22]; the “ImageNet Strong” baselines, including $\mathtt{GradNorm}$ [17], $\mathtt{NECO}$ [1], $\mathtt{React}$ [38], $\mathtt{Dice}$ [37], $\mathtt{ASH}$ [8], $\mathtt{Scale}$ [47]. See Appendix 9 for details of the baselines.
Performance Table 1(a) shows that $\mathtt{NCI}$ consistently ranks top-three across benchmarks, whereas baselines exhibit greater variability. To assess overall performance, we averaged AUROC across benchmarks, which are of a similar range. Table 1(b) highlights that NCI improves overall performance compared to strong baselines on individual benchmarks. Further, $\mathtt{NCI}$ is as efficient as $\mathtt{MSP}$, as shown in Table 1(b) [^2] Running time of $\mathtt{KNN}$ on ImageNet are copied from Table 4 in [39]., which enhances efficiency compared to strong baselines. This aligns with the analysis in Section 3 and Appendix 9. We highlight the following pairs of comparison:
  • $\mathtt{NCI}$ v.s. $\mathtt{NCI}$ w/o filter: On the CIFAR-10 classifier, strong ID clustering allows our method to rank top-3 without filtering. Conversely, on the ImageNet ResNet-50, weaker ID clustering (see Appendix 13) makes norm-based filtering crucial for reducing generalization discrepancy. Complete Neural Collapse occurs on neither model while $\mathtt{NCI}$ remains effective.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{KNN}$: Compared to $\mathtt{KNN}$, $\mathtt{NCI}$ significantly reduces the latency (Table 1(b)). Notably, without filtering, our hyperparameter-free score outperforms KNN with tuned parameters on most benchmarks (Table 1(a), Table 2(a) & Table 9), highlighting the benefit of using class-specific information.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{ASH}$ / $\mathtt{Scale}$: Compared to both, $\mathtt{NCI}$ delivers competitive performance on ImageNet and significantly improves CIFAR-10, enhancing overall performance ( Table 1(b)). Also, $\mathtt{ASH}$ and $\mathtt{Scale}$ introduce in a small delay on the ImageNet benchmark due to activation sorting, with larger activation dimensions likely widening the latency gap on larger models.
@@
  • $\mathtt{NCI}$ v.s. $\mathtt{NECO}$: $\mathtt{NECO}$ [1] is motivated by Neural Collapse. Like $\mathtt{NCI}$ with filtering, $\mathtt{NECO}$ uses max-logit and incorporates distance to the origin. However, $\mathtt{NECO}$ exclusively analyzes features, requiring expensive matrix multiplication and leading to higher inference latency (Table 1(b)). Conversely, $\mathtt{NCI}$ explores the interplay between features and the classification head, integrating class-specific information to improve both efficiency and effectiveness.
4.2 Mitigating Discrepancies across Architectures
@@
Next, we study two transformer-based models: ViT B/16 [9] and Swin-v2 [26], both finetuned on ImageNet, achieving an accuracy of 81.14% and 82.94% respectively. We follow the setup of the OpenOOD ImageNet Benchmark in Section 4.1. Based on validation results, we set the filter strength $\alpha$ of the $L1$ norm to $10^{-3}$ for both classifiers. In Table 2, we observe strong baselines suffer on Swin v2, echoing the observations in [1]. Conversely, our $\mathtt{NCI}$, even without filtering, improves baseline performance on Swin v2. Filtering further enhances the performance, leading to improved overall performance (Table 2(b)).
Table 3: NCI improves the overall performance, averaged across Table 1(a), Table 2(a) & Table 9.
We further aggregate in Table 2 with experiments on ResNet (Table 1) and DenseNet (Table 9) and report the average AUROC in Table 3. $\mathtt{NCI}$ significantly boosts the overall performance.
@@ 4.3 Ablation on the Filtering Effect
In Table 4, we assess different orders of $p$-norm as the filtering term, compared to the $L1$ norm used so far. To ensure a fair comparison, we report the best performance from the filter strengths $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. The rest of the setup follows the ImageNet benchmarks in Section 4.1. As shown in Table  4, filtering with $L1$ norm achieves the best performance across OOD datasets, aligning with prior observations [17, 33]. Meanwhile, we observe that in rare scenarios, e.g., a ResNet-18 on CIFAR-10, the $\mathtt{L1}$ norm cannot effectively characterize OOD’s proximity to the origin, leading to no extra performance gain compared to simply thresholding on $\mathtt{pScore}$. In these cases, our algorithm benefits from its ability to automatically select a low filter strength based on validation results, effectively disregarding the filtering term.
Table 4: Ablation on filtering norm on ImageNet OpenOOD Benchmark with ResNet-50 backbone. AUROC score is reported (higher is better). Bold denotes the best. Filtering with $\mathtt{L1}$ norm outperforms alternative choice of norms across OOD datasets.
We also test the sensitivity of $\mathtt{NCI}$ to filtering strength $\alpha$ in Table 5. As shown on the ImageNet ResNet50 benchmark, performance remains stable for $\alpha$ values within the same scale. Given this insensitivity, we select hyperparameters from four scales $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$ without extensive finetuning in this work.
Table 5: Sensitivity of $\mathtt{NCI}$ to filtering strength. Average AUROC on ImageNet ResNet-50 Benchmark reported. Performance remains stable within the same scale.
We further apply $\mathtt{L1}$-norm based filtering to $\mathtt{KNN}$ to see if this perspective can mitigate the discrepancy of clustering-based methods in general. In Table 6 333 Note that we report our run of KNN here to ensure a fair evaluation of the filtering effect. Our results are very similar to the OpenOOD results reported in Table 1(a), with only marginal differences. , we report the the best performance of $\mathtt{KNN}$ from filter strengths $\{10^{-4},10^{-3},10^{-2},10^{-1}\}$. We observe a significant performance gain from adding the filter, which further validates our understanding of ID clustering landscape from Neural Collapse. Note that our method outperforms the standalone $L1$ norm as well as $\mathtt{KNN}$, before and after filtering.